Find the area of an equilateral triangle with each side 2 using dθrdr NOT rdrdθ
Answer #1

Area of equilateral triangle is

A=a^2\sqrt{3}/4

we have a=2

A=2^2.\sqrt{3}/4

A=1.73 unit^2

Area of equilateral triangle with each side 2 is 1. 73 unit square.